OpenStudy (anonymous):

fractions: 3/X = 39/182 Find X

jhonyy9 (jhonyy9):

39x=3*182

jhonyy9 (jhonyy9):

divide both sides by 3

OpenStudy (anonymous):

divide the nominator on both sides. 39 / 3 = 13 Now divide the denominator with 13 182 / 13 = 14 39 / 182 = 3 / 14 X = 14

jhonyy9 (jhonyy9):

13x=182

OpenStudy (anonymous):

x=(3*182)/39 =14

OpenStudy (anonymous):

Thanks guys!

OpenStudy (anonymous):

X= 14